OpenStudy (broskishelleh):

A sandbag was thrown downward from a building. The function f(t) = -16t2 - 64t + 80 shows the height f(t), in feet, of the sandbag after t seconds: Part A: Factor the function f(t) and use the factors to interpret the meaning of the x-intercept of the function. (4 points) Part B: Complete the square of the expression for f(x) to determine the vertex of the graph of f(x). Would this be a maximum or minimum on the graph? (4 points) Part C: Use your answer in part B to determine the axis of symmetry for f(x)? (2 points)

OpenStudy (broskishelleh):

Please halp

Nnesha (nnesha):

given expression is f(t) where is f(x) or is it a typo ?

OpenStudy (broskishelleh):

F(t) is F(x), just replace t with x if you want

Nnesha (nnesha):

alright let's factor the right side i have few mints sorry

OpenStudy (broskishelleh):

It's fine

Nnesha (nnesha):

\[\huge\rm f(t)= \color{ReD}{-16t^2-64t+80}\] first take out the common factor

OpenStudy (broskishelleh):

The common factor is 8

Nnesha (nnesha):

well we need GREATEST common factor

OpenStudy (broskishelleh):

Oh, 16 then my mistake

Nnesha (nnesha):

yes right and bec leading coefficient is negative we can take out the negative 16

Nnesha (nnesha):

and sorry gtg :(

OpenStudy (broskishelleh):

Alright that leaves us with \[f(t) = -4 + 5\]

OpenStudy (broskishelleh):

Okay bye

Nnesha (nnesha):

aw didn't get the help :( sorry

Nnesha (nnesha):

alright when we find the common denominator we should divide all terms by the common factor \[\large\rm \color{blue}{\frac{ -16t^2 }{ -16 }-\frac{ 64t }{ -16}+\frac{ 80 }{ -16} }\]\[\large \rm f(t)=-16(\color{blue}{t^2+4t-5})\] now factor t^2+4t-5 there are more than 3 ways to factor o the quadratic equation you can use AC method(easy) find two number when you multiply them you should get product of AC (a=leading coefficient , c=constant term) and when you add or subtract them you should get middle term(b)

Nnesha (nnesha):

remember t^2 is same as 1t^2 so in this question leading coefficient is 1